OpenStudy (anonymous):

evalute the integral (sin^3 (x)/cosx)dx

hartnn (hartnn):

hi @jasper21dude write sin^3 x = sin^2 x (sin x dx) and then plug in u= cos x

hartnn (hartnn):

try that substitution and tell me what u get ?

OpenStudy (shubhamsrg):

you may write sin^3(x) /cos(x) as tan^3 (x) sec^2(x) now a substitution of tanx = t should help .

OpenStudy (shubhamsrg):

I mean both methods are fine. Second one perhaps a bit shorter ?

hartnn (hartnn):

difference of one or 2 steps is not actually a difference, both methods are equivalent :)

OpenStudy (shubhamsrg):

hmm..right.

OpenStudy (raden):

tan^3 (x) sec^2(x) = sin^3(x)/cos^3(x) * 1/cos^2(x) = sin^3(x)/cos^5(x) @shubhamsrg

OpenStudy (raden):

it means not equal (sin^3 (x)/cosx)

hartnn (hartnn):

oohhh....hemight have thought, tan^3 x cos^2 x ...which doesn't help i guess.....

OpenStudy (raden):

my idea is (sin^3 (x)/cosx) = sin^2 (x) sinx/cosx) = (cos^2 x - 1) sinx/cosx = sinxcosx - tanx = 1/2 sin2x - tanx so, it becomes int 1/2 sin2x dx - int tanx dx

OpenStudy (raden):

damn... thts wrong too :)

hartnn (hartnn):

yes, just that sin^2x = 1- cos^2 x

OpenStudy (raden):

well, corrction :) (sin^3 (x)/cosx) = sin^2 (x) sinx/cosx) = (1 - cos^2(x) ) sinx/cosx = tanx - sinxcos = tanx - 1/2sin2x so, it is int tan(x) dx - 1/2 int sin2x dx =

OpenStudy (raden):

that should be easy for you @jasper21dude

hartnn (hartnn):

alternatively, after = (1 - cos^2(x) ) sinx/cosx if we put u= cos x du = -sin x dx we get, integral of (u^2-1)/u which is equally easy to integrate (just don't forget to substitute back u=cos x)

OpenStudy (raden):

wow, interesting :)

OpenStudy (shubhamsrg):

oops! :P sorry..